# Pan-LUT: Efficient Pan-sharpening via Learnable Look-Up Tables

### v1

@@ Abstract
Recently, deep learning-based pan-sharpening algorithms have achieved notable advancements over traditional methods. However, many deep learning-based approaches incur substantial computational overhead during inference, especially with high-resolution images. This excessive computational demand limits the applicability of these methods in real-world scenarios, particularly in the absence of dedicated computing devices such as GPUs and TPUs. To address these challenges, we propose Pan-LUT, a novel learnable look-up table (LUT) framework for pan-sharpening that strikes a balance between performance and computational efficiency for high-resolution remote sensing images. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Our proposed method contains fewer than 300K parameters and processes a 8K resolution image in under 1 ms using a single NVIDIA GeForce RTX 2080 Ti GPU, demonstrating significantly faster performance compared to other methods. Experiments reveal that Pan-LUT efficiently processes large remote sensing images in a lightweight manner, bridging the gap to real-world applications. Furthermore, our model surpasses SOTA methods in full-resolution scenes under real-world conditions, highlighting its effectiveness and efficiency.
* Equal contributions  $\dagger$ Corresponding author.
@@ 1 Introduction
High-resolution multispectral (HRMS) images are widely used in applications such as military operations, environmental monitoring, and mapping. However, due to the limitations of physical sensors, these images are challenging to obtain. Pan-sharpening addresses this issue by fusing high-resolution panchromatic (PAN) images with low-resolution multispectral (LRMS) images, producing high-quality HRMS images through complementary integration [47] [55][61].
Figure 1: Comparisons of performance and computational efficiency. The PSNR value is evaluated on the Worldview-III dataset. Meanwhile, the inference time is measured on images with a resolution of ($2048\times 2048$) by employing a single NVIDIA GeForce RTX 2080 Ti GPU. Our method outperforms traditional approaches by a significant margin of 7 dB, while maintaining a speed comparable to that of traditional methods. Furthermore, our approach even surpasses some DNN-based approaches, outperforming Pan-GAN by more than 0.3 dB while being 145.82 times faster. Note that (4K, 8K) refer to models processing images at resolutions of $4096\times 4096$ and $8192\times 8192$, respectively.
Recently, numerous pan-sharpening methods have been proposed, which can be generally categorized into two main groups: traditional methods and deep learning-based methods. Traditional pan-sharpening methods, such as component substitution (CS) [2] [3] [11], multi-resolution analysis (MRA) [7] [20], and variational optimization (VO) [10] [34], often struggle to restore precise spatial and spectral details in HRMS images. This limitation arises from their heavy reliance on manually designed features and insufficient modeling of prior knowledge. In contrast, deep learning-based pan-sharpening methods have demonstrated exceptional fusion capabilities due to the powerful feature extraction ability of deep neural networks (DNNs). Masi et al. [32] built the Pan-sharpening Neural Network (PNN) model, which first applied CNN to pan-sharpening field, achieving a significant improvement over classical methods. Following this, researchers have explored more complicated and deeper networks to further promote the performance of pan-sharpening [12] [39] [33] [40] [56]. Besides, transformer-based models have garnered attention in pan-sharpening for their remarkable ability in long-range modeling, such as [53] [57] [59]. Nevertheless, many existing approaches suffer from significant computational overhead and memory costs during inference, particularly when dealing with high-resolution images. These excessive demands limit the practical deployment of these methods, especially in environments lacking dedicated computing resources such as GPUs and TPUs. As shown in  Figure 1, very few methods are capable of processing 4K-resolution images under limited computational resources. Even among the deep learning-based methods that can handle 2K-resolution images, their inference speed is often suboptimal. In real-world applications, remote sensing images typically have much higher resolutions, and these methods may encounter significant challenges in operating efficiently under such conditions. Furthermore, simply increasing the depth of a network does not guarantee improved performance, as deeper networks become more difficult to train and may suffer from overfitting due to an excessive number of redundant parameters. To this end, our work aims to overcome these challenges by introducing a more efficient and lightweight pan-sharpening approach that delivers high performance and computational efficiency for high-resolution remote sensing images.
The lookup table (LUT), as a vital constituent in image signal processors (ISPs) [19], owes its significance to several advantageous characteristics, such as runtime efficiency and hardware-friendly nature. They offer a promising data structure for efficiently performing specific transformations by replacing expensive runtime computations with inexpensive array caching and indexing. However, the internal parameters of LUTs are usually preset by experts, making it difficult to optimize them for specific images, which limits their effectiveness.
@@
To overcome the aforementioned challenges, we propose a novel learnable LUT framework, called Pan-LUT, which achieves a good balance between performance and computational efficiency in pan-sharpening. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Our method consists of fewer than 300K parameters and can processes a 8K resolution image in under 1 ms using a single NVIDIA GeForce RTX 2080 Ti GPU. Furthermore, our approach outperforms traditional methods by 7 dB, while maintaining a speed comparable to that of conventional techniques, demonstrating superior speed and efficiency compared to existing methods.
In summary, the contributions are as follows:
@@
  • We present Pan-LUT, a novel learnable LUT framework that does not incorporate any network structure. This framework is designed to achieve a strong balance between performance and computational efficiency in pan-sharpening high-resolution remote sensing images.
@@
  • To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT).
@@
Figure 2: The overall framework of our proposed Pan-LUT. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT).
2 Related Works
@@ 2.1 Look-Up Table
Lookup Table (LUT) is a widely-used data structure in Image Signal Processors (ISPs) [5] [8] [18], which maps a set of indices to corresponding output values, allowing efficient retrieval without the need for repeated computations. LUTs are particularly useful for functions of multiple variables, as they store precomputed outputs for all possible input combinations. For example, in a 1D LUT, a single input index is mapped to an output value, often using linear interpolation for indices that fall between pre-stored values. More complex LUTs, such as 3D LUTs, use three independent input variables, which may require advanced interpolation methods like trilinear or tetrahedral interpolation. Due to its portability, various LUT based solutions have been proposed for image enhancement [6] [23] [26, 27]. For instance, Zeng et al. [52] and Wang et al. [38] propose image-adaptive 3D LUTs for efficient single-image enhancement. These approaches rely on a network weight predictor to fuse different 3D LUTs, which may pose a limitation on platforms under resource-constrained conditions. Additionally, LUT-based methods have been explored in the area of super-resolution [17] [22] [30] [28]. SRLUT [17] trains a deep super-resolution (SR) network with a restricted receptive field and then cache the output values from the learned SR network in LUTs. However, issues such as performance degradation arise when large patches are cached in LUTs, prompting the development of strategies like MuLUT [22], which introduces multiple LUT variants and a fine-tuning strategy to improve performance. To further enhance the functionality of LUTs, architectures like SPLUT [30] and RCLUT [28] have been proposed. SPLUT processes different image information separately using multiple LUTs, while RCLUT introduces a plugin module to improve LUT-based models with minimal additional computational cost.
@@ 2.2 Traditional Pan-sharpening Methods
Traditional fusion techniques encompass component substitution (CS), multi-resolution analysis (MRA), and variational optimization (VO). CS methods, such as IHS [2], Brovey [11], and PCA [3], utilize spatial details from high-resolution panchromatic (PAN) images to replace corresponding details in low-resolution multispectral (LRMS) images, often causing spectral distortion due to the incomplete incorporation of spectral information. MRA techniques, including DWT [20] and ATWT [7], apply multi-resolution decomposition to merge PAN and LRMS images, which helps to reduce spectral distortion. VO methods, such as Bayesian [10] and Total Variation [34], formulate the fusion process as an optimization problem by iteratively minimizing the loss function. While VO methods show promising results, they encounter challenges in optimizing model design and loss functions. Although these approaches have yielded certain improvements, their performance remains constrained by the inadequate modeling, which restricts further advancements in pan-sharpening accuracy and quality.
@@
Brovey [11] 35.8646 0.9216 0.0403 1.8238 37.7974 0.9026 0.0218 1.3720 22.5060 0.5466 0.1159 8.2331 0.20 0.21 0.23
IHS [2] 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616 0.20 0.21 0.23
SFIM [29] 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730 0.26 0.29 0.33
GS [21] 35.6376 0.9176 0.0423 1.8774 37.2260 0.9034 0.0309 1.6736 22.5608 0.5470 0.1217 8.2433 0.68 0.73 0.76
PNN [32] 40.7550 0.9624 0.0259 1.0646 43.1208 0.9704 0.0172 0.8528 29.9418 0.9121 0.0824 3.3206 4.21 12.81 54.59
MSDCNN [50] 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884 16.45 49.82 OOM
Pan-GAN [31] 39.1025 0.9562 0.0303 1.2954 41.4468 0.9661 0.0205 1.0593 28.4959 0.8897 0.0998 3.9067 14.55 40.83 OOM
GPPNN [44] 41.1622 0.9684 0.0244 1.0315 44.2145 0.9815 0.0137 0.7361 30.1785 0.9175 0.0776 3.2593 17.59 43.60 OOM
SFDI [60] 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5462 30.5971 0.9236 0.0741 3.0798 20.26 65.37 OOM
UCGAN [58] 40.0545 0.9553 0.0275 1.1734 42.3634 0.9557 0.0194 0.9480 28.6705 0.8851 0.0990 3.8696 17.26 52.06 OOM
PanFlow [48] 41.8584 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 3.1142 18.27 54.19 OOM
PSCINN [37] 41.8520 0.9703 0.0223 0.9407 47.1100 0.9878 0.0107 0.5612 30.5599 0.9230 0.0748 3.1033 28.31 60.74 OOM
Pan-Mamba [13] 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1551 0.9299 0.0702 2.8942 42.31 87.74 OOM
TA-DiffHQP [41] 42.1255 0.9752 0.0211 0.9023 47.7716 0.9900 0.0101 0.5378 31.3369 0.9302 0.0737 2.6369 481.27 998.58 OOM
Pan-LUT (Ours) 39.8362 0.9555 0.0286 1.1695 42.6559 0.9642 0.0189 0.9044 28.8213 0.8936 0.0935 3.7520 0.27 0.28 0.30
2.3 Deep Learning-based Methods
Deep learning-based methods have emerged as the dominant approach for pan-sharpening in recent years. The PNN [32] model, inspired by SRCNN [9], is the first to introduce CNNs into this domain, surpassing traditional methods. Models like PanNet [49] and MSDCNN [50] further enhance performance by leveraging residual connections and multi-scale convolutions, effectively capturing high-frequency details and supporting a wide range of remote sensing applications. Since then, more complex CNN-based architectures [4] [14] [15] [62] [41] have been proposed in this field to improve pan-sharpening’s mapping ability. Models like GPPNN [44], MMNet [46], and ARFNet [45], which enhance interpretability through deep unfolding techniques and accelerate model convergence. Other approaches, including SFDI [60] and MSDDN [50], utilize Fourier transforms to capture high-frequency features. Transformer-based architectures, such as INN-former [59], Panformer [57] and DRFormer [53] combine CNNs and Transformers to capture both local and global features. Instead of learning a deterministic mapping, generative models such as UCGAN [58], PanFlow [48] and PSCINN [37] generate a distribution of possible outputs for the given inputs. Additionally, some networks [25] [54] [42] [24] focus on improving the generalization capability of CNN-based fusion models by learning generalized features. Despite their promising results, these advanced methods come with high computational costs, limiting their practical applicability.
3 Method
3.1 Framework
The overall framework of our proposed Pan-LUT is illustrated in Figure 2. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT) and adaptive aggregation look-up table (AALUT). Specifically, given the PAN image ($P\in R^{H\times W\times 1}$) and the upsampled MS image ($MS\in R^{H\times W\times C}$), they are first concatenated to form $PM\in R^{H\times W\times(C+1)}$, which is then fed into PGLUT for channel-wise spectral mapping, producing the output $V_{c}\in R^{H\times W\times C}$. SDLUT, using only the PAN image as input, generates local spatial details $V_{s}\in R^{H\times W\times C}$. To further enhance spatial representation, AALUT takes $V_{s}\in R^{H\times W\times C}$ as input to adaptively learn local contexts, resulting in $V_{as}\in R^{H\times W\times C}$. By summing up $V_{c}$, and $V_{as}$, we can obtain the final HRMS result:
$\begin{gathered}V_{c}=PGLUT(PM),\\ V_{s}=SDLUT(P),V_{as}=AALUT(V_{s}),\\ HRMS=V_{c}+V_{as}.\\ \end{gathered}$ (1)
3.2 Spectral Transformation
To preserve the rich spectral information of the MS image, we propose the PAN-guided look-up table (PGLUT), which leverages the PAN image as guidance to finely control the spectral transformation. Specifically, PGLUT is represented as a 5-dimensional matrix containing $N^{5}$ elements, where $N$ denotes the number of bins in each dimension. Each element corresponds to a sampling point, defining a set of indexed input pixels $\{I_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$ and their corresponding output pixels $\{O_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$. Here, $I\in\{pa,r,g,b,nir\}$ represents the pixel values from the PAN and MS images, while $O\in\{R,G,B,NIR\}$ denotes the corresponding cached output pixels.
Since the distribution of LUT elements is discrete in space, the output value cannot be directly retrieved from the LUT. For an input value $\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image. PGLUT first performs a lookup operation to locate the corresponding input pixel in the LUT:
$\begin{gathered}x=\frac{pa_{(w,h)}^{I}}{V_{max}}\cdot N,y=\frac{r_{(w,h)}^{I}}{V_{max}}\cdot N,z=\frac{g_{(w,h)}^{I}}{V_{max}}\cdot N,\\ s=\frac{b_{(w,h)}^{I}}{V_{max}}\cdot N,e=\frac{nir_{(w,h)}^{I}}{V_{max}}\cdot N,\\ \end{gathered}$ (2)

### v2

@@ Abstract
Recently, deep learning-based pan-sharpening algorithms have achieved notable advancements over traditional methods. However, deep learning-based methods incur substantial computational overhead during inference, especially with large images. This excessive computational demand limits the applicability of these methods in real-world scenarios, particularly in the absence of dedicated computing devices such as GPUs and TPUs. To address these challenges, we propose Pan-LUT, a novel learnable look-up table (LUT) framework for pan-sharpening that strikes a balance between performance and computational efficiency for large remote sensing images. Our method makes it possible to process $\times15K$ remote sensing images on a 24GB GPU. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details, we introduce the spatial details look-up table (SDLUT). Furthermore, to adaptively aggregate channel information for generating high-resolution multispectral images, we design an adaptive output look-up table (AOLUT). Our model contains fewer than 700K parameters and processes a $\times9K$ image in under 1 ms using one RTX 2080 Ti GPU, demonstrating significantly faster performance compared to other methods. Experiments reveal that Pan-LUT efficiently processes large remote sensing images in a lightweight manner, bridging the gap to real-world applications. Furthermore, our model surpasses SOTA methods in full-resolution scenes under real-world conditions, highlighting its effectiveness and efficiency. The source code is available at https://github.com/CZhongnan/Pan-LUT.
@@ 1 Introduction
High-resolution multispectral (HRMS) images are widely used in applications such as military operations, environmental monitoring, and mapping. However, due to the limitations of physical sensors, these images are challenging to obtain. Pan-sharpening addresses this issue by fusing high-resolution panchromatic (PAN) images with low-resolution multispectral (LRMS) images, producing high-quality HRMS images through complementary integration zheng2023deep zhou2023pan wang2023learning. Recently, numerous pan-sharpening methods have been proposed, which can be generally categorized into two main groups: traditional methods and deep learning-based methods. Traditional methods, such as component substitution (CS) ihs pca brovey, multi-resolution analysis (MRA) ATWT DWT, and variational optimization (VO) Bayesian TV, often struggle to restore precise spatial or spectral details in HRMS images. In contrast, deep learning-based pan-sharpening methods have demonstrated exceptional fusion capabilities due to the powerful feature extraction ability of deep neural networks (DNNs) lin2024pair lin2024difftv lin2025jarvisir lin2024aglldiff. Masi et al. pnn built the Pan-sharpening Neural Network (PNN) model, which first applied CNN to pan-sharpening field, achieving a significant improvement over traditional methods.
Figure 1: Comparisons of computational efficiency. Our method can process $\times9K$ and $\times15K$ images on GPUs with 11GB and 24GB memory, respectively. Meanwhile, we observe that (a) DNN-based methods are highly sensitive to the image size, and (b) in the absence of a GPU, they require a considerable amount of time to process images. In the CPU inference time experiments, all methods were conducted on a workstation equipped with an Intel(R) Xeon(R) Gold 6226R CPU.
Following this, researchers have explored more complicated and deeper networks to further promote the performance of pan-sharpening py_2023_ICCV lin2023domain meng2024progressive zhong2025ssdiff. However, they overlook a critical practical issue: the need for real-time processing of large remote sensing images in real-world applications. As illustrated in Figure 1, we observe two major limitations in DNN-based approaches: (1) they are highly sensitive to the size of the input images, and (2) they rely heavily on dedicated computing devices such as GPUs and TPUs. Increasing GPU memory does not significantly improve the image size these methods can handle and several of these methods demand a significant amount of time to process images in CPU-only environments. And we will demonstrate in the experiments sections that, even with GPU acceleration, most methods still fail to process large remote sensing images in real time. In practical applications, remote sensing images typically exhibit even higher resolutions, posing additional challenges to existing methods in terms of efficiency and scalability. Moreover, simply increasing network depth does not necessarily lead to better performance, as deeper models are harder to train and often suffer from overfitting due to redundant parameters.
@@
To overcome the aforementioned challenges, we propose a novel learnable Look-Up Table (LUT) framework, called Pan-LUT, which achieves a good balance between performance and computational efficiency in pan-sharpening. Specifically, we replace complex DNN operations with learnable LUTs to enable lightweight deployment in practical applications. To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details, we introduce the spatial details look-up table (SDLUT). To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT). The Pan-LUT consists of fewer than 700K parameters and can processes $\times9K$ images in under 1 ms using a single RTX 2080 Ti GPU. Furthermore, our approach outperforms traditional methods by 7 dB, while maintaining a speed comparable to that of conventional techniques, demonstrating superior speed and efficiency compared to existing methods. Our contributions can be summarized as follows:
@@
  • We present Pan-LUT, a novel learnable LUT framework that does not incorporate any network structure. This framework is designed to achieve a strong balance between performance and computational efficiency in pan-sharpening high-resolution remote sensing images. Our method makes it possible to process $\times15K$ remote sensing images on one 24GB GPU.
@@
  • To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping. To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT). To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT).
@@
2 Related Work
@@ 2.1 Look-Up Table
Look-up Tables (LUT) are particularly useful for functions of multiple variables, as they store precomputed outputs for all possible input combinations. For example, in a 1D LUT, a single input index is mapped to an output value, often using linear interpolation for indices that fall between pre-stored values. More complex LUTs, such as 3D LUTs, use three independent input variables, which may require advanced interpolation methods like trilinear or tetrahedral interpolation. Due to its portability, various LUT based solutions have been proposed for image enhancement conde2024nilut li2023fastllve lin2024unsupervised salut 3dlut. For instance, Zeng et al. 3dlut and Wang et al. salut propose image-adaptive 3D LUTs for efficient single-image enhancement. These approaches rely on a network weight predictor to fuse different 3D LUTs, which may pose a limitation on platforms under resource-constrained conditions. Additionally, LUT-based methods have been explored in the area of super-resolution srlut mulut splut rclut. SRLUT srlut trains a deep super-resolution (SR) network with a restricted receptive field and then cache the output values from the learned SR network in LUTs. However, issues such as performance degradation arise when large patches are cached in LUTs, prompting the development of strategies like MuLUT mulut, which introduces multiple LUT variants and a fine-tuning strategy to improve performance. To further enhance the functionality of LUTs, architectures like SPLUT splut and RCLUT rclut have been proposed. SPLUT processes different image information separately using multiple LUTs, while RCLUT introduces a plugin module to improve LUT-based models with minimal additional computational cost.
Figure 2: The overall framework of our proposed Pan-LUT. PGLUT is a spectral transformation LUT designed to extract spectral information, SDLUT is a spatial detail transformation LUT for capturing texture features, and AOLUT is an adaptive output LUT used to aggregate channel information.
@@ 2.2 Traditional Pan-sharpening Methods
Traditional fusion techniques encompass component substitution (CS), multi-resolution analysis (MRA), and variational optimization (VO). CS methods, such as IHS ihs, Brovey brovey, and PCA pca, utilize spatial details from high-resolution panchromatic (PAN) images to replace corresponding details in low-resolution multispectral (LRMS) images, which can lead to spectral distortion due to the incomplete incorporation of spectral information. MRA techniques, including DWT DWT and ATWT ATWT, apply multi-resolution decomposition to merge PAN and LRMS images, which enables better preservation of spectral information and reduces spectral distortion. VO methods, such as Bayesian Bayesian and Total Variation TV, formulate the fusion process as an optimization problem by iteratively minimizing the loss function. While VO methods show promising results, they encounter challenges in optimizing model design and loss functions. Although these approaches have yielded certain improvements, their performance remains constrained by the inadequate modeling, which restricts further advancements in pan-sharpening accuracy and quality.
2.3 Deep Learning-based Methods
Deep learning-based methods have emerged as the dominant approach for pan-sharpening in recent years wang2024cross he2024frequency wang2025towards. The PNN pnn model, inspired by SRCNN srcnn, is the first to introduce CNNs into this domain, surpassing traditional methods. Models like PanNet pannet and MSDCNN msdcnn further enhance performance by leveraging residual connections and multi-scale convolutions, effectively capturing high-frequency details and supporting a wide range of remote sensing applications. Since then, more complex CNN-based architectures chen2024novel hou2023bidomain zhu2023probability have been proposed in this field to improve the mapping ability of pan-sharpening. Models like GPPNN gppnn, MMNet mmnet, and ARFNet arfnet, which enhance interpretability through deep unfolding techniques and accelerate model convergence. Spatial adaptive convolution methods, such as LAGConv jin2022lagconv and CANConv duan2024canconv, can adaptively generate different convolution ker nel parameters based on various spatial locations, enabling them to accommodate different spatial regions. Other approaches, including SFDI sfiin and MSDDN msdcnn, utilize Fourier transforms to capture high-frequency features. Transformer-based architectures, such as INN-former inn, Panformer zhou2022panformer and DRFormer zhang2023drformer combine CNNs and Transformers to capture both local and global features. Instead of learning a deterministic mapping, generative models such as UCGAN ucgan, PanFlow  yang2023panflownet and PSCINN wang2024panpsci generate a distribution of possible outputs for the given inputs. In addition, several studies have investigated lightweight pan-sharpening methods, including SpanConv chen2022spanconv and LGPConv zhao2023lgpconv. However, we show that current lightweight methods are only lightweight in the context of low-resolution images. Despite their promising results, these advanced methods come with high computational costs, limiting their practical applicability.
3 Method
3.1 Framework
The overall framework of our proposed Pan-LUT is illustrated in Figure 2, which consists of three specifically designed LUTs. (1) To finely control the spectral transformation, we devise the PAN-guided look-up table (PGLUT) for channel-wise spectral mapping, which incorporates a PAN-guided indexing strategy and a pentalinear interpolation technique. (2) To effectively capture fine-grained spatial details and adaptively learn local contexts, we introduce the spatial details look-up table (SDLUT), which incorporates a rotation-enhanced indexing strategy and a quadrilinear interpolation technique. (3) To further enable adaptive channel aggregation for high-resolution multispectral image generation, we design the adaptive output look-up table (AOLUT), which incorporates a PAN-guided indexing strategy and a pentalinear interpolation technique. Specifically, given the PAN image ($P\in\mathbb{R}^{H\times W\times 1}$) and the upsampled MS image ($MS\in\mathbb{R}^{H\times W\times C}$), they are concatenated into $PM\in\mathbb{R}^{H\times W\times(C+1)}$ and passed through the PGLUT for channel-wise spectral mapping, producing the output $V_{pg}\in\mathbb{R}^{H\times W\times(C+1)}$. SDLUT then takes $V_{pg}$ as input to generate local spatial details, yielding $V_{sd}\in\mathbb{R}^{H\times W\times(C+1)}$. Finally, AOLUT takes $V_{sd}$ as input to generate the final $HRMS\in R^{H\times W\times C}$ result:
[…]
3.2 Spectral Transformation
To preserve the rich spectral information of the MS image, we propose the PAN-guided look-up table (PGLUT), which leverages the PAN image as guidance to finely control the spectral transformation. Specifically, PGLUT is represented as a 5-dimensional matrix containing $N^{5}$ elements, where $N$ denotes the number of bins per dimension. Each element corresponds to a sampling point, defining a set of indexed input pixels $\{I_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$ and their corresponding output pixels $\{O_{(i,j,k,m,n)}\}_{i,j,k,m,n=0,...,N-1}$. Here, $I\in\{pa,r,g,b,nir\}$ represents the pixel values from the PAN and MS images, while $O\in\{R,G,B,NIR\}$ denotes the corresponding cached output pixels. Since the LUT elements are discretely distributed in space, the output value cannot be directly retrieved from the LUT. For an input value $\{pa_{(w,h)}^{I},r_{(w,h)}^{I},g_{(w,h)}^{I},b_{(w,h)}^{I},nir_{(w,h)}^{I}\}$, where $(w,h)$ denotes the spatial position of a pixel in the image.
PAN-guided indexing strategy. As illustrated in Figure 2, we introduce an indexing strategy for precise spectral transformation, referred to as the PAN-guided indexing strategy. In a MS image, pixels from different spatial locations may have identical values (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j}$, where $i\neq j$). The LUT maps these identical inputs to the same output. The PAN-guided indexing strategy provides a more flexible indexing mechanism for the LUT. Specifically, it additionally considers the pixels at corresponding spatial positions in the PAN image (e.g., $r_{i}=r_{j},g_{i}=g_{j},b_{i}=b_{j},nir_{i}=nir_{j},pa_{i}\neq pa_{j}$, where $i\neq j$), thereby achieving finer-grained mapping. Specifically, PGLUT first performs a lookup operation to locate the corresponding input pixel in the LUT:
[…]
where $\lfloor\cdot\rfloor$ denotes the floor function. $\{\mathbf{d}_{l}\}_{l=x,y,z,s,e}$ represents the offset of the input index $(x,y,z,s,e)$ relative to the defined sampling point $(i,j,k,m,n)$, e.g., $\mathbf{d}_{x}={x}-{i}$.
Pentalinear Interpolation. After locating 32 adjacent points, an appropriate interpolation technique is applied to these sampled values to generate the output:
$O_{(x,y,z,s,e)}=PInterpolation(LUT[L],{\{\mathbf{d}_{l}\}}),$ (4)
where $PInterpolation(\cdot)$ denotes the pentalinear interpolation. More details about PGLUT and the pentalinear interpolation can be found in Section A.2.
3.3 Spatial Details Transformation
PGLUT is essentially a channel-wise 5D LUT that operates globally, which limits its ability to capture local spatial information. To effectively capture fine-grained spatial details and adaptively learn local contexts, we propose the Spatial Details Lookup Table (SDLUT).
Rotation-indexing strategy. As illustrated in Figure 2, given a pixel $p_{(w,h)}$, SDLUT processes this pixel along with its neighboring pixels as input. During the training phase, we employ a Rotation-indexing strategy to further expand the receptive field, which can be formulated as:
[…]
where $f_{SDLUT}(\cdot)$ denotes the lookup and interpolation process in the LUT retrieval. More details about SDLUT and the quadrilinear interpolation can be found in Section A.3.
@@ 3.3 Spatial Details Transformation
Brovey 35.8646 0.9216 0.0403 1.8238 37.7974 0.9026 0.0218 1.3720 22.5060 0.5466 0.1159 8.2331 - 0.28 0.33
IHS 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616 - 0.23 0.26
SFIM 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730 - 0.32 0.47
GS 35.6376 0.9176 0.0423 1.8774 37.2260 0.9034 0.0309 1.6736 22.5608 0.5470 0.1217 8.2433 - 0.75 0.87
PNN 40.7550 0.9624 0.0259 1.0646 43.1208 0.9704 0.0172 0.8528 29.9418 0.9121 0.0824 3.3206 0.0689 12.81 54.59
PanNet 40.8176 0.9626 0.0257 1.0557 43.0659 0.9685 0.0178 0.8577 29.6840 0.9072 0.0851 3.4263 0.0688 29.52 OOM
MSDCNN 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884 0.2390 49.82 OOM
Pan-GAN 39.1025 0.9562 0.0303 1.2954 41.4468 0.9661 0.0205 1.0593 28.4959 0.8897 0.0998 3.9067 0.0915 40.83 OOM
GPPNN 41.1622 0.9684 0.0244 1.0315 44.2145 0.9815 0.0137 0.7361 30.1785 0.9175 0.0776 3.2593 0.1198 43.60 OOM
SFDI 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5462 30.5971 0.9236 0.0741 3.0798 0.0871 65.37 OOM
UCGAN 40.0545 0.9553 0.0275 1.1734 42.3634 0.9557 0.0194 0.9480 28.6705 0.8851 0.0990 3.8696 0.2109 52.06 OOM
PanFlow 41.8584 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 3.1142 0.0873 54.19 OOM
PSCINN 41.8520 0.9703 0.0223 0.9407 47.1100 0.9878 0.0107 0.5612 30.5599 0.9230 0.0748 3.1033 3.3209 60.74 OOM
Pan-Mamba 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1551 0.9299 0.0702 2.8942 0.1827 87.74 OOM
TA-DiffHQP 42.1255 0.9752 0.0211 0.9023 47.7716 0.9900 0.0101 0.5378 31.3369 0.9302 0.0737 2.6369 2.6000 998.58 OOM
Pan-LUT (Ours) 40.8555 0.9633 0.0254 1.0339 43.7466 0.9726 0.0169 0.8027 29.7376 0.9106 0.0815 3.3934 0.6626 0.38 0.54
3.4 Adaptive Output
For the feature channel pixels from the SDLUT $(V_{(w,h)}^{1},V_{(w,h)}^{2},V_{(w,h)}^{3},V_{(w,h)}^{4},V_{(w,h)}^{5})$, AOLUT adaptively aggregates channel information to generate high-resolution multispectral channel pixels $\{R_{(w,h)}^{O},G_{(w,h)}^{O},B_{(w,h)}^{O},NIR_{(w,h)}^{O}\}$. Pixel-level Transformation can be formulated as: